OpenStudy (anonymous):

Traffic signs are regulated by the Manual on Uniform Traffic Control Devices (MUTCD). The perimeter of some traffic sign must be 96 inches. Also, its length is 10 inches longer than its width. Find the dimensions of this sign. The width of the sign is: The length of the sign is:

OpenStudy (anonymous):

what is the shape? rectangular i guess

OpenStudy (anonymous):

Sorry yes I think it's rectangular

OpenStudy (anonymous):

uhm. x=19 and y=29

OpenStudy (anonymous):

w=x l=x+10 p=96 2(x+10+x)=96 2x+10=48 2x=38 x=19

OpenStudy (anonymous):

w=19 l=29

OpenStudy (anonymous):

Thank you both so much.

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

uw